The following data describes the position of an object along the x -axis as a function of time. Plot the data and find the instantaneous velocity of the object, at a-5.0s, b-16s,c-23s


1
Expert's answer
2021-08-02T14:33:17-0400

Gives

Graph value between position and time

Find velocity

"t=5sec\\t=16sec\\t=23sec"


"t=5sec"

"V_{t=5sec}=\frac{11.3-7.8}{5}=0.018m/sec"

"+x" direction

"t=16sec\\V_{t=16sec}=\frac{20-20}{16}=0m/sec"

t=23sec


"V_{t=23sec}=\frac{13.5-16.2}{23}=-0.013m/sec"

"-x" direction